# Qingdao 266237, China

### v1

@@ Appendix 0.A Leakage Analysis of [GNT23]
Gao et al. [GNT23] claimed that their MPSU protocol is secure in the presence of arbitrary colluding participants. However, we find that their protocol fails to achieve this security, and also requires the same non-collusion assumption as LG protocol. First, we give a brief review of their protocol.
@@
Apart from MKR-PKE, the protocol of [GNT23] utilizes three primary ingredients: 1) The multi-point OPRF [PRTY19, CM20, RS21, RR22], where the sender $\mathcal{S}$ inputs a PRF key $k$, and the receiver $\mathcal{R}$ inputs $n$ elements $\{x_{1},\cdots,x_{n}\}$ and receives the corresponding PRF evaluations $\{F_{k}(x_{1}),\cdots,F_{k}(x_{n})\}$. 2) The conditional oblivious pseudorandom function (cOPRF), an extension they develop on the multi-point OPRF [^17] In their paper, the definition of cOPRF is the single-point version. For ease of illustration, we use the multi-point version., where the sender $\mathcal{S}$ additionally inputs a set $Y$. If $x_{i}\notin Y$, $\mathcal{R}$ receives $F_{k}(x_{i})$, else $\mathcal{R}$ receives a random value sampled by $\mathcal{S}$ [^18] In the latest version of [GNT23], they have made a slight adjustment to the functionality of cOPRF. However, this adjustment does not change their insecurity. Their latest protocol is still vulnerable to our attack. 3) The membership Oblivious Transfer (mOT), where $\mathcal{S}$ inputs an element $y$ and two messages $u_{0},u_{1}$, while $\mathcal{R}$ inputs a set $X$ and receives $u$, one of $u_{0},u_{1}$. If $y\notin X$, $u=u_{0}$, else $u=u_{1}$.
To illustrate the insecurity of their protocol, we consider a four-party case where the leader $P_{1}$ possesses a set $X_{1}$ of items while the remaining parties $P_{2},P_{3},P_{4}$, each possesses a single item $X_{i}$ = $\{x_{i}\}$ ($i=2,3,4$). We assume that $x_{2},x_{3},x_{4}\notin X_{1}$ and $x_{2}\neq x_{3}=x_{4}$ [^19] In their paper, to illustrate the high-level idea of the protocol, they give a four-party case as well, assuming that $x_{2}=x_{3}\notin X_{1}$ and $x_{4}\in X_{1}$, see [GNT23]..
[…]
Figure 14: Illustration of the mPSU protocol in [GNT23] for $m=4$ parties. $P_{1}$ possesses a set $X_{1}=\{x_{1}^{1},\cdots,x_{1}^{n}\}$ of $n$ items while $P_{i}$ has input set of only one item $x_{i}$ for $i=2,3,4$. We assume that $x_{2},x_{3},x_{4}\notin X_{1}$ and $x_{2}\neq x_{3}=x_{4}$. $P_{1}$ possesses two set $S,E$ and updates them in each round. In the $i$-th round, $P_{1}$ and $P_{i}$ invoke mpOPRF and mOT in sequence. $F_{k}(X_{1})$ denotes the set of PRF values on $X_{1}$, i.e., $\{F_{k}(x_{1}^{1}),\cdots,F_{k}(x_{1}^{n})\}$. $P_{i}$ invokes cOPRF and mOT with $P_{i+1},\cdots,P_{m}$ concurrently. At the end of $i$-th round, $S$ contains the multi-key PRF values on $X_{1}\cup\cdots\cup X_{i+1}$ using keys $k_{2},\cdots,k_{i+1}$, $E$ contains the MKR-PKE ciphertexts of $X_{1}\cup\cdots\cup X_{i+1}\setminus X_{1}$. Finally, all parties cooperate in a Shuffle&Decrypt protocol to decrypt $E$ in a random permutation. We denote $F_{k_{3}}(F_{k_{2}}(\cdot))$ as $F_{k_{2},k_{3}}(\cdot)$, more precisely, it is in the form of $H(k_{3},H(k_{2},\cdot\|1)\|1)$ in the paper. We emphasize that the ciphertext $\mathsf{Enc}(pk,x_{3})$ in Round 1 is different from that in Round 2. The latter is the rerandomized ciphertext of the former. We write them using the same denotation as this does not affect our illustration of the the attack. Changes to sets $S$ and $E$ at the end of each round are marked in blue font.
We demonstrate the execution process of their MPSU in Figure 14 pertaining to the scenario we describe above. In light of space constraints, Figure 14 only covers Round 1 and Round 2 and it omits some invocations of cOPRF and mOT since the potential information leakage already occurs during Round 2 when $P_{1}$ and $P_{3}$ invoke the mOT. To elaborate, $P_{1}$ receives $F_{k_{2}}(x_{2})$ from the invocation of mOT with $P_{2}$ in Round 1 and $F_{k_{2},k_{3}}(x_{3})$ from the invocation of mOT with $P_{3}$ in Round 2. If $P_{1}$ and $P_{3}$ collude, $P_{1}$ can compute $F_{k_{2},k_{3}}(x_{2})$ using $P_{3}$’s key $k_{3}$ and learn the knowledge of $F_{k_{2},k_{3}}(x_{2})\neq F_{k_{2},k_{3}}(x_{3})$, which indicates $x_{2}\neq x_{3}$. Recall that $P_{1}$ is the leader to obtain the union $X_{1}\cup\{x_{2},x_{3}=x_{4}\}$ in the end. Therefore, the coalition of $P_{1}$ and $P_{3}$ can deduce that the unique element in the union, not present in their joint sets $X_{1}\cup\{x_{3}\}$ (i.e., $x_{2}$), belongs to $P_{2}$. Accordingly, they know that $P_{4}$ possesses the duplicate element with either $P_{1}$ or $P_{3}$. Note that in the ideal world, the coalition of $P_{1}$ and $P_{3}$ should not acquire knowledge of the source party of the foreign element in the union or the party possessing the duplicate element. Therefore, their protocol at least requires the same assumption as LG protocol, assuming that the leader $P_{1}$ is not in collusion with any other parties.
@@ Appendix 0.B Public-Key Encryption
Appendix 0.C Construction of Multi-Party Secret-Shared Shuffle
The functionality of share correlation generated in the offline phase is depicted in Figure 15. The online details of multi-party secret-shared shuffle in [EB22] are given in Figure 16.
  • Functionality. On input $\pi_{i}:[n]\to[n]$ from each $P_{i}$ ($1\leq i\leq m$), sample $\vec{a}_{i}^{\prime},\vec{b}_{i}\leftarrow(\{0,1\}^{l})^{n}$ for $11\leq i\leq m-$ 1, and $\vec{a}_{i},\vec{\Delta}_{m}\leftarrow(\{0,1\}^{l})^{n}$ for $2\leq i\leq m$, which satisfy
[…]
    Give $\vec{a}_{1}^{\prime},\vec{b}_{1}$ to $P_{1}$, $\vec{a}_{i}^{\prime},\vec{a}_{i},\vec{b}_{i}$ to $P_{i}$ ($1<i<m$), and $\vec{a}_{m},\vec{\Delta}_{m}$ to $P_{m}$.
Figure 15: Share Correlation Functionality $\mathcal{F}_{\mathsf{sc}}$
  • Parameters: $m$ parties $P_{1},\cdots,P_{m}$. Ideal functionality $\mathcal{F}_{\mathsf{sc}}$ in Figure 15. The dimension of vector $n$. The item length $l$.
  • Inputs: Each party $P_{i}$ has input $\vec{x}_{i}={(x_{i}^{1},\cdots,x_{i}^{n})}$.
  • Protocol:
    1. 1.
      Each party $P_{i}$ samples a random permutation $\pi_{i}:[n]\to[n]$ and invokes $\mathcal{F}_{\mathsf{sc}}$ with input $\pi_{i}$. $P_{1}$ receives $\vec{a}_{1}^{\prime},\vec{b}_{1}\in(\{0,1\}^{l})^{n}$; For $2\leq j<m$, $P_{j}$ receives $\vec{a}_{j}^{\prime},\vec{a}_{j},\vec{b}_{j}\in(\{0,1\}^{l})^{n}$; $P_{m}$ receives $\vec{a}_{m},\vec{\Delta}_{m}\in(\{0,1\}^{l})^{n}$.
    2. 2.
      For $2\leq j\leq m$, $P_{j}$ computes $\vec{c}_{j}=\vec{x}_{j}\oplus\vec{a}_{j}$ and sends $\vec{c}_{j}$ to $P_{1}$.
    3. 3.
      $P_{1}$ computes $\vec{c^{\prime}}_{1}=\pi_{1}(\bigoplus_{j=2}^{m}\vec{c}_{j}\oplus\vec{x}_{1})\oplus\vec{a}_{1}^{\prime}$ and send it to $P_{2}$. $P_{1}$ outputs $\vec{b}_{1}$.
    4. 4.
      For $2\leq j<m$, $P_{j}$ computes $\vec{c^{\prime}}_{j}=\pi_{j}(\vec{c^{\prime}}_{j-1})\oplus\vec{a}_{j}^{\prime}$ and send it to $P_{j+1}$. $P_{j}$ outputs $\vec{b}_{j}$.
    5. 5.
      $P_{m}$ outputs $\pi_{m}(\vec{c^{\prime}}_{m-1})\oplus\vec{\Delta}_{m}$.
Figure 16: Multi-Party Secret-Shared Shuffle Protocol $\Pi_{\mathsf{ms}}$
Appendix 0.D Construction of MKR-PKE from ElGamal
The MKR-PKE is instantiated using ElGamal encryption as follows:
@@ Appendix 0.D Construction of MKR-PKE from ElGamal
  • The randomized encryption algorithm $\mathsf{Enc}$ takes as input a public key $pk$ and a plaintext message $x\in\mathbb{G}$, samples $r\leftarrow\mathbb{Z}_{q}$, and outputs a ciphertext $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2})=(g^{r},x\cdot pk^{r})$.
@@
Appendix 0.E Missing Security Proofs
0.E.1 The Proof of Theorem 5.1
  • Parameters: $m$ parties $P_{1},\cdots,P_{m}$. The message length $l$.
  • Inputs: $P_{1}$ has input $b_{1}\in\{0,1\}$ and $P_{m}$ has input $b_{m}\in\{0,1\}$.
  • Protocol:
    1. 1.
      $P_{1}$ and $P_{m}$ invoke $\mathcal{F}_{\mathsf{rot}}$ twice: First, $P_{1}$ acts as receiver with input $b_{1}$ and $P_{m}$ acts as sender. $P_{1}$ receives $r_{m,1}^{b_{1}}\in\{0,1\}^{l}$. $P_{m}$ receives $r_{m,1}^{0},r_{m,1}^{1}\in\{0,1\}^{l}$; Second, $P_{m}$ acts as receiver with input $b_{m}$ and $P_{1}$ acts as sender. $P_{m}$ receives $r_{1,m}^{b_{m}}\in\{0,1\}^{l}$. $P_{1}$ receives $r_{1,m}^{0},r_{1,m}^{1}\in\{0,1\}^{l}$.
    2. 2.
      For $1<j<m$: $P_{1}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{1}$ acts as receiver with input $b_{1}$ and $P_{j}$ as sender without input. $P_{1}$ receives $r_{j,1}^{b_{1}}\in\{0,1\}^{l}$. $P_{j}$ receives $r_{j,1}^{0},r_{j,1}^{1}\in\{0,1\}^{l}$.
    3. 3.
      For $1<j<m$: $P_{m}$ and $P_{j}$ invoke $\mathcal{F}_{\mathsf{rot}}$ where $P_{m}$ acts as receiver with input $b_{m}$ and $P_{j}$ as sender. $P_{m}$ receives $r_{j,m}^{b_{m}}\in\{0,1\}^{l}$. $P_{j}$ receives $r_{j,m}^{0},r_{j,m}^{1}\in\{0,1\}^{l}$.
    4. 4.
      For $1<j<m$: $P_{j}$ samples $\Delta_{j}\leftarrow\{0,1\}^{l}$ and computes $r_{j}=r_{j,1}^{0}\oplus r_{j,m}^{0}$. $P_{j}$ sends $u_{j,1}=\Delta_{j}\oplus r_{j,1}^{0}\oplus r_{j,1}^{1}$ to $P_{1}$, and $u_{j,m}=\Delta_{j}\oplus r_{j,m}^{0}\oplus r_{j,m}^{1}$ to $P_{m}$, then outputs $(r_{j},\Delta_{j})$.
    5. 5.
      $P_{1}$ computes $\Delta_{1}=r_{1,m}^{0}\oplus r_{1,m}^{1}$ and $r_{1}=\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1}$ ($\cdot$ denotes bitwise-AND between the repetition code of $b_{1}$ and $u_{j,1}$, which are both strings of length $l$. Similarly hereinafter), then outputs $(r_{1},\Delta_{1})$. $P_{m}$ computes $\Delta_{m}=r_{m,1}^{0}\oplus r_{m,1}^{1}$ and $r_{m}=\bigoplus_{j=2}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus r_{1,m}^{b_{m}}\oplus b_{m}\cdot\Delta_{m}$, then outputs $(r_{m},\Delta_{m})$.
In order to facilitate the proof, we assume that $\text{ch}_{0}=1$ and $\text{ch}_{1}=m$, and $P_{1}$ and $P_{m}$ provide inputs $b_{1}\in\{0,1\}$ and $b_{m}\in\{0,1\}$ respectively. We turn to proving the correctness and security of the protocol in Figure 17. Note that when $\text{ch}_{0}$ and $\text{ch}_{1}$ are assigned different values, the proof is essentially the same.
@@ Appendix 0.E Missing Security Proofs
    $r_{1}\oplus(\bigoplus_{j=2}^{m-1}r_{j})\oplus r_{m}=(\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1})$
    $\oplus(r_{j1}^{0}\oplus r_{j,m}^{0})\oplus(\bigoplus_{j=2}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus r_{m,1}^{b_{m}}\oplus b_{m}\cdot\Delta_{m})$ (1)
@@
    $r_{j,1}^{b_{1}}=r_{j,1}^{0}\oplus b_{1}\cdot(r_{j,1}^{0}\oplus r_{j,1}^{1}),r_{j,m}^{b_{m}}=r_{j,m}^{0}\oplus b_{m}\cdot(r_{j,m}^{0}\oplus r_{j,m}^{1}),$ (3)
    $r_{m,1}^{b_{1}}=r_{m,1}^{0}\oplus b_{1}\cdot(r_{j,1}^{0}\oplus r_{j,1}^{1})=r_{m,1}^{0}\oplus b_{1}\cdot\Delta_{1},$ (4)
    $r_{1,m}^{b_{m}}=r_{1,m}^{0}\oplus b_{m}\cdot(r_{1,m}^{0}\oplus r_{1,m}^{1})=r_{1,m}^{0}\oplus b_{m}\cdot\Delta_{m}$ (5)
@@
    $r_{1}\oplus(\bigoplus_{j=2}^{m-1}r_{j})\oplus r_{m}$
    $=(b_{1}\oplus b_{m})\cdot\Delta_{1}\oplus(\bigoplus_{j=2}^{m-1}(b_{1}\oplus b_{m})\cdot\Delta_{j})\oplus(b_{1}\oplus b_{m})\cdot\Delta_{m}$
    $=\bigoplus_{i=1}^{m}(b_{1}\oplus b_{m})\cdot\Delta_{i}$
    Then we can summarize that if $b_{1}\oplus b_{m}=0$, $r_{1}=\bigoplus_{j=2}^{m}r_{j}$, else $r_{1}=\Delta_{1}\oplus(\bigoplus_{j=2}^{m}(r_{j}\oplus\Delta_{j}))$. This is exactly the functionality $\mathcal{F}_{\mathsf{mss\text{-}rot}}$.
@@
    To elaborate, in the case that $P_{1}\notin\mathcal{C}orr$ and $P_{m}\notin\mathcal{C}orr$, simulator receives all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$ and needs to emulate each $P_{c}$’s view, including its private $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$. The simulator for corrupted $P_{c}$ runs the protocol honestly except that it sets $P_{c}$’s random tape to be the output $\Delta_{c}$, and simulates uniform outputs from $\mathcal{F}_{\mathsf{rot}}$ under the constraint that $r_{c,1}^{0}\oplus r_{c,m}^{0}=r_{c}$. Clearly, the joint distribution of all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
@@
    In the case that $P_{1}\in\mathcal{C}orr$ or $P_{m}\in\mathcal{C}orr$, since the protocol is symmetric with respect to the roles of $P_{1}$ and $P_{m}$, we focus on the case of corrupted $P_{1}$. The simulator receives $b_{1}$ in addition to all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$. For $P_{1}$, its view consists of its private $\Delta_{1}$, outputs $r_{m,1}^{b_{1}}$, $(r_{1,m}^{0},r_{1,m}^{1})$, $\{r_{j,1}^{b_{1}}\}_{1<j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,1}\}_{1<j<m}$ from $P_{j}$. For each $P_{c}(c\neq 1)$, its view consists of its private $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$.
@@
    The simulator sets each corrupted party $P_{c}$’s random tape to be its output $\Delta_{c}$. Then for $P_{1}$’s view, it runs the protocol honestly except that it simulates uniform outputs $r_{m,1}^{b_{1}}$, $(r_{1,m}^{0},r_{1,m}^{1})$, $r_{i,1}^{b_{1}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniformly random messages $u_{i,1}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m-1}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus r_{m,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{1}=r_{1}$, where $P_{i}\in\mathcal{H}$. For other corrupted parties’ view, it runs the protocol honestly except that it sets $r_{c,m}^{0}=r_{c,1}^{0}\oplus r_{c}$ and simulates uniform output $r_{c,m}^{1}$ from $\mathcal{F}_{\mathsf{rot}}$.
@@
    In the real execution, $P_{1}$ receives $u_{i,1}=\Delta_{i}\oplus r_{i,1}^{0}\oplus r_{i,1}^{1}$. From the definition of ROT functionality, $r_{i,1}^{0}$ (when $b_{1}=0$) or $r_{i,1}^{1}$(when $b_{1}=0$) is uniform and independent of $P_{1}$’s view. Therefore, $u_{i,1}$ is uniformly at random from the perspective of $P_{1}$. Clearly, the joint distribution of $b_{1}$ and all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
@@
    In the case that $P_{1}\in\mathcal{C}orr$ and $P_{m}\in\mathcal{C}orr$, the simulator receives $b_{1},b_{m}$ and all outputs $(r_{c},\Delta_{c})$ of $P_{c}\in\mathcal{C}orr$. For $P_{1}$, its view consists of its private $\Delta_{1}$, outputs $r_{m,1}^{b_{1}}$, $(r_{1,m}^{0},r_{1,m}^{1})$, $\{r_{j,1}^{b_{1}}\}_{1<j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,1}\}_{1<j<m}$ from $P_{j}$. For each $P_{c}(c\neq 1,c\neq m)$, its view consists of its private $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$. For $P_{m}$, its view consists of its private $\Delta_{m}$, outputs $r_{1,m}^{b_{m}}$, $(r_{m,1}^{0},r_{m,1}^{1})$, $\{r_{j,1}^{b_{m}}\}_{1<j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,m}\}_{1<j<m}$ from $P_{j}$.

### v4

@@ Appendix 0.A Leakage Analysis of [GNT23]
    The MPSU protocol in [GNT23] is claimed to be secure in the presence of arbitrary colluding participants. However, our analysis would suggest that the protocol fails to achieve this security, and also requires the non-collusion assumption as LG. First, we give a brief review of the protocol.
@@
    Apart from MKR-PKE, their protocol utilizes two new ingredients: 1) The conditional oblivious pseudorandom function (cOPRF), an extension they develop on the OPRF, where the sender $\mathcal{S}$ additionally inputs a set $Y$. If $x\notin Y$, $\mathcal{R}$ receives $F_{k}(x)$, else $\mathcal{R}$ receives a random value sampled by $\mathcal{S}$. 2) The membership Oblivious Transfer (mOT), where $\mathcal{S}$ inputs an element $y$ and two messages $u_{0},u_{1}$, while $\mathcal{R}$ inputs a set $X$ and receives $u$, one of $u_{0},u_{1}$. If $y\in X$, $u=u_{0}$, else $u=u_{1}$.
    To illustrate insecurity of their protocol, we consider a three-party case where $P_{1}$ and $P_{3}$ each possess a single item $X_{1}=\{x_{1}\}$ and $X_{3}=\{x_{3}\}$ respectively, while $P_{2}$ possesses a set $X_{2}$. We assume that $x_{1}=x_{3}$. According to the protocol (cf. Figure 8 in their paper), in step 3.(a), $P_{1}$ and $P_{2}$ invoke the OPRF where $P_{1}$ acts as $\mathcal{R}$ inputting $x_{1}$ and $P_{2}$ acts as $\mathcal{S}$ inputting its PRF key $k_{2}$. $P_{1}$ receives the PRF value $F_{k_{2}}(x_{1})$. Meanwhile, in step 3.(c), $P_{2}$ and $P_{3}$ invoke the cOPRF where $P_{3}$ acts as $\mathcal{R}$ inputting $x_{3}$, and $P_{2}$ acts as $\mathcal{S}$ inputting its PRF key $k_{2}$ and the set $X_{2}$. $P_{3}$ receives the output $w$ from the cOPRF. By the definition of cOPRF functionality, if $x_{3}\notin X_{2}$, $w=F_{k_{2}}(x_{3})$, otherwise $w$ is a random value.
    If $P_{1}$ and $P_{3}$ collude, they can distinguish the cases where $x_{3}\in X_{2}$ and $x_{3}\notin X_{2}$ by comparing $P_{1}$’s output $F_{k_{2}}(x_{1})$ from the OPRF and $P_{3}$’s output $w$ from the cOPRF for equality. To elaborate, we recall that $x_{1}=x_{3}$, so $F_{k_{2}}(x_{1})=F_{k_{2}}(x_{3})$. If $F_{k_{2}}(x_{1})=w$, it implies that $P_{3}$ receives $F_{k_{2}}(x_{3})$ from the cOPRF, so the coalition learns that $x_{3}\notin X_{2}$; On the contrary, if $F_{k_{2}}(x_{1})\neq w$, it implies that $P_{3}$’s output from the cOPRF is not $F_{k_{2}}(x_{3})$, so it is a random value, then the coalition learns that $x_{3}\in X_{2}$. More generally, as long as $P_{1}$ and $P_{3}$ collude, they can identify whether each element $x\in X_{1}\cap X_{3}$ belongs to $X_{2}$ or not, by comparing the PRF value $F_{k}(x)$ from the OPRF between $P_{1}$ and $P_{2}$ and the cPRF value (whose condition depends on $x\in X_{2}$ or not) from the cOPRF between $P_{2}$ and $P_{3}$. This acquired knowledge is information leakage in MPSU. Therefore, the non-collusion assumption is required.
@@ Appendix 0.B Public-Key Encryption
    Appendix 0.C Construction of MKR-PKE from ElGamal
    MKR-PKE is instantiated with ElGamal encryption below:
@@ Appendix 0.C Construction of MKR-PKE from ElGamal
    • The randomized encryption algorithm $\mathsf{Enc}$ takes as input a public key $pk$ and a plaintext message $x\in\mathbb{G}$, samples $r\leftarrow\mathbb{Z}_{q}$, and outputs $\mathsf{ct}=(\mathsf{ct}_{1},\mathsf{ct}_{2})=(g^{r},x\cdot pk^{r})$.
@@
    Appendix 0.D Missing Security Proofs
    0.D.1 The Proof of Theorem 5.1
    We consider the case that $\text{ch}_{0}=1$ and $\text{ch}_{1}=m$, and $J=\{1,\cdots,m\}$. Let $P_{1}$ and $P_{m}$ input $b_{1}\in\{0,1\}$ and $b_{m}\in\{0,1\}$, and $P_{i}$ ($1\leq i\leq m$) input $\Delta_{i}$ respectively. We turn to proving the correctness and security of the protocol in Figure 10 in this particular case. Note that in the different cases, the proof is essentially the same.
@@ Appendix 0.D Missing Security Proofs
      $r_{1}=\bigoplus_{j=2}^{m}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus b_{1}\cdot\Delta_{1},$ (1)
      $r_{i}=r_{i,1}^{0}\oplus r_{i,m}^{0},1<i<m,$ (2)
      $r_{m}=\bigoplus_{j=1}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot u_{j,m})\oplus r_{m,1}^{0}\oplus b_{m}\cdot\Delta_{m},$ (3)
@@
      $r_{j,1}^{b_{1}}=r_{j,1}^{0}\oplus b_{1}\cdot(r_{j,1}^{0}\oplus r_{j,1}^{1}),$ (5)
      $r_{j,m}^{b_{m}}=r_{j,m}^{0}\oplus b_{m}\cdot(r_{j,m}^{0}\oplus r_{j,m}^{1}),$ (6)
@@
      $r_{1}=\bigoplus_{j=2}^{m}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot\Delta_{j})\oplus r_{1,m}^{0}\oplus b_{1}\cdot\Delta_{1},$ (7)
      $r_{m}=\bigoplus_{j=1}^{m-1}(r_{j,m}^{b_{m}}\oplus b_{m}\cdot\Delta_{j})\oplus r_{m,1}^{0}\oplus b_{m}\cdot\Delta_{m},$ (8)
      Substitute Equation 278 into $r_{1}\oplus(\bigoplus_{j=2}^{m-1}r_{j})\oplus r_{m}$, we have $r_{1}\oplus(\bigoplus_{j=2}^{m-1}r_{j})\oplus r_{m}=\bigoplus_{i=1}^{m}(b_{1}\oplus b_{m})\cdot\Delta_{i}$. Then we can summarize that if $b_{1}\oplus b_{m}=0$, $r_{1}=\bigoplus_{j=2}^{m}r_{j}$, else $r_{1}=\Delta_{1}\oplus(\bigoplus_{j=2}^{m}(r_{j}\oplus\Delta_{j}))$. This is exactly the functionality $\mathcal{F}_{\mathsf{mss\text{-}rot}}$.
@@
      To elaborate, in the case that $P_{1}\notin\mathcal{C}orr$ and $P_{m}\notin\mathcal{C}orr$, simulator receives all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$ and needs to emulate each $P_{c}$’s view, including its private input $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$. The simulator for corrupted $P_{c}$ runs the protocol honestly except that it simulates uniform outputs from $\mathcal{F}_{\mathsf{rot}}$ under the constraint that $r_{c,1}^{0}\oplus r_{c,m}^{0}=r_{c}$. Clearly, the joint distribution of all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
@@
      In the case that $P_{1}\in\mathcal{C}orr$ or $P_{m}\in\mathcal{C}orr$, since the protocol is symmetric with respect to the roles of $P_{1}$ and $P_{m}$, we focus on the case of corrupted $P_{1}$. The simulator receives all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$. For $P_{1}$, its view consists of the choice bit $b_{1}$, its private input $\Delta_{1}$, outputs $(r_{1,m}^{0},r_{1,m}^{1})$, $\{r_{j,1}^{b_{1}}\}_{1<j\leq m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,1}\}_{1<j\leq m}$ from $P_{j}$. For each $P_{c}(c\neq 1)$, its view consists of its private input $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$.
@@
      For $P_{1}$’s view, the simulator runs the protocol honestly except that it simulates uniform outputs $(r_{1,m}^{0},r_{1,m}^{1})$, $r_{i,1}^{b_{1}}$ from $\mathcal{F}_{\mathsf{rot}}$ and uniformly random messages $u_{i,1}$ from $P_{i}$ under the constraint $\bigoplus_{j=2}^{m}(r_{j,1}^{b_{1}}\oplus b_{1}\cdot u_{j,1})\oplus r_{1,m}^{0}\oplus b_{1}\cdot\Delta_{1}=r_{1}$, where $P_{i}\in\mathcal{H}$. For other corrupted parties’ view, it runs the protocol honestly except that it sets $r_{c,m}^{0}=r_{c,1}^{0}\oplus r_{c}$ and simulates uniform output $r_{c,m}^{1}$ from $\mathcal{F}_{\mathsf{rot}}$.
@@
      In the real execution, $P_{1}$ receives $u_{i,1}=\Delta_{i}\oplus r_{i,1}^{0}\oplus r_{i,1}^{1}$. From the definition of ROT functionality, $r_{i,1}^{0}$ (when $b_{1}=0$) or $r_{i,1}^{1}$(when $b_{1}=0$) is uniform and independent of $P_{1}$’s view. Therefore, $u_{i,1}$ is uniformly at random from the perspective of $P_{1}$. Clearly, the joint distribution of all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$, along with their view emulated by simulator, is indistinguishable from that in the real process.
@@
      In the case that $P_{1}\in\mathcal{C}orr$ and $P_{m}\in\mathcal{C}orr$, the simulator receives all outputs $r_{c}$ of $P_{c}\in\mathcal{C}orr$. For $P_{1}$, its view consists of the choice bit $b_{1}$, its private input $\Delta_{1}$, outputs $(r_{1,m}^{0},r_{1,m}^{1})$, $\{r_{j,1}^{b_{1}}\}_{1<j\leq m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,1}\}_{1<j\leq m}$ from $P_{j}$. For each $P_{c}(c\neq 1,c\neq m)$, its view consists of its private input $\Delta_{c}$, outputs $(r_{c,1}^{0},r_{c,1}^{1})$ and $(r_{c,m}^{0},r_{c,m}^{1})$ from $\mathcal{F}_{\mathsf{rot}}$. For $P_{m}$, its view consists of its private input $\Delta_{m}$, outputs $(r_{m,1}^{0},r_{m,1}^{1})$, $\{r_{j,m}^{b_{m}}\}_{1\leq j<m}$ from $\mathcal{F}_{\mathsf{rot}}$ and protocol messages $\{u_{j,m}\}_{1\leq j<m}$ from $P_{j}$.